# Endmember Extraction from Hyperspectral Images Using Self-Dictionary Approach with Linear Programming

### v2

@@ VI-A Computational Efficiency
We chose ten equally spaced points between 00 and $1$ and set them as noise intensity levels $\nu$. One dataset contained ten HSI matrices $A=W[I,\bar{H}]+V$ of size $50\times n$ with $r=10$ satisfying $\nu=\|V\|_{1}$ for each $\nu$. We set $n$ from $500$ to $2500$ in $500$ increments and, thus, constructed five datasets in total.
@@
We compared RCE-SR and -DIR with CPLEX-SR and -DIR, since the Hottopixx methods in [16, 19, 28] uses CPLEX and solves the Hottopixx models directly. Figure 1 summarizes the average elapsed time for each dataset. We can see from the figure that RCE is faster than CPLEX for $n\geq 1000$. In particular, RCE-SR is significantly faster than the other methods and its average elapsed time increases more slowly. RCE-DIR is faster than CPLEX-DIR, but its average elapsed time increases more rapidly than RCE-SR. The results suggest that RCE-SR should be used for the implementation of Hottopixx when dealing with large matrices.
Figure 1: Average elapsed time of four methods for five datasets.
VI-B Endmember Extraction Performance
Next, we examined the endmember extraction performance of EEHT on semi-real HSI datasets. The datasets were constructed from the Jasper Ridge and Samson HSI datasets. We extracted subimages in the manner described in Section V of [33]:
@@ VI-B Endmember Extraction Performance
The top of Figure 5 in Appendix F displays the RGB images of these datasets.
The endmember signatures in these regions were already identified in [33]. Using them as reference signatures, we constructed $W,H$ and $V$ as follows. Let $A^{\mathrm{real}}$ be the HSI matrix for a dataset, and $\bm{w}^{\mathrm{ref}}_{1},\ldots,\bm{w}^{\mathrm{ref}}_{r}$ be the reference signature vectors. Perform the following procedure:
@@
    Compute the optimal solution $X$ of problem (4) by letting $(A,\mathcal{I})=(A^{\mathrm{real}},\mathcal{J})$ in the problem. Set $X(:,\mathcal{J})=I$ and then $H=X$.
@@
We constructed two datasets using the $W,H$ and $V$ obtained above: dataset 1 from Jasper Ridge and dataset 2 from Samson. Each dataset contained $20$ HSI matrices that were generated as follows: choose $20$ equally spaced points between 00 and $1$ as noise intensity levels $\nu$; and construct an HSI matrix $A=WH+(\nu/\|V\|_{1})\cdot V$ for each $\nu$. The HSI matrices $A$ were of size $198\times 10000$ with $4$ endmembers for dataset 1 and of size $156\times 9025$ with $3$ endmembers for dataset 2. Note that $A=A^{\mathrm{real}}$ holds if $\nu=\|V\|_{1}$, where $\|V\|_{1}\approx 0.61$ for dataset 1 and $\|V\|_{1}\approx 0.15$ for dataset 2.
@@
Figure 2: Average MRSA value (×$\times 10^{2}$) of EEHT and existing methods on dataset 1 (top) and dataset 2 (bottom).
@@
For a fair comparison with EEHT, we included the SVD-based dimensionality reduction, described in Section V, in MERIT: i.e., first compute the top-$r$ truncated SVD $A_{r}=U_{r}\Sigma_{r}V_{r}^{\top}$ of the input matrix $A$; then run MERIT on $A^{\prime}=\Sigma_{r}V_{r}^{\top}$.
We ran EEHT and the six existing methods on datasets 1 and 2, and evaluated their performance by using the average MRSA given in (11). Regarding MERIT, it is necessary to fix the parameter $\lambda$ of problem (6) and the parameter $\mu$ of the function $\phi_{\mu}$ shown in (7) before it. In particular, the choice of $\lambda$ affects its endmember extraction performance. We set $\mu$ to the default value of $10^{-5}$, as suggested by the authors of [30]. We then set $\lambda$ to $10^{-6},10^{-5},...,10^{3}$ and ran MERIT for each of these 10 values. For each $\lambda$ value, we calculated the mean of the average MRSA values across all intensity levels $\nu$. For dataset 1 (resp., dataset 2), the means were at their minimum when $\lambda=10^{2}$ (resp., $10^{-1}$). Therefore, we report the results of MERIT with $\lambda=10^{2}$ for dataset 1 and $\lambda=10^{-1}$ for dataset 2 in Figure 2.
@@
Figure 2 plots the average MRSA values of the methods for datasets 1 and 2. We can see from the figure that, in most cases, EEHT-C achieved better average MRSA values than EEHT-A, -B, and the five greedy methods. This result suggests that using EEHT-C is preferable to EEHT-A and -B. EEHT-C achieved worse average MRSA values than MERIT for dataset 1, while it performed better than MERIT for dataset 2. For the $20$ HSI matrices included in each dataset, the computation time of EEHT-A, -B and -C was approximately 30 minutes on average.
@@
VI-C Hyperspectral Unmixing of Urban HSI
@@
Finally, we conducted an experimental study on hyperspectral unmixing of the Urban HSI dataset. The image was taken over Copperas Cove, Texas, USA by the HYDICE sensor. It consists of $307\times 307$ pixels with $210$ bands from 400 nm to 2500 nm. Following the procedure in [33, 32], we removed dirty bands from the image. The resulting image had 162 clean bands. The experimental studies in [33, 32] showed that there are mainly $4$-$6$ endmembers in the image. Following the settings used in [20], our experiments set the number of endmembers to $6$: Asphalt, Grass, Tree, Roof 1, Roof 2 and Soil. These six endmember signatures were identified in [32]. We used them as reference signatures. The bottom of Figure 5 in Appendix F displays the RGB image of the Urban dataset.
@@
We did not achieve satisfactory results for Urban by applying the methods that were tested on the semi-real HSI datasets described in Section VI-B. The estimated endmember signatures were not sufficiently close to the reference ones. Consequently, we conducted data-specific preprocessing aimed at enhancing the endmember extraction performance of these methods. The preprocessing was based on a reasonable assumption that there are a number of pixels close to each pure pixel in an HSI. Indeed, this assumption holds in the case of the Urban dataset: see Table I that summarizes the number of pixels under an MRSA value of $0.05$ relative to the reference signatures of the endmembers. Here, we shall say that pixels are isolated if there are almost no pixels close to them. Under this assumption, even if we remove isolated pixels, there would be little effect on the number of pure pixels. It is possible that some of the isolated pixels contain large amounts of noise, which would worsen the endmember extraction performance of the methods.
@@
TABLE I: Number of pixels under an MRSA value of $0.05$ relative to the reference signatures of endmembers for Urban.
@@
To find isolated pixels, we introduced the neighborhood density of each pixel, defined by
@@
$\mathrm{NH}(i;\phi)=\left\{j\in\mathcal{N}\ \middle|\ \mathrm{MRSA}(\bm{a}_{i},\bm{a}_{j})\leq\phi\right\},$
and the total number $n$ of pixels in an HSI. For Urban, we examined the neighborhood density $\rho$ while varying $\phi$. Figure 3 displays a histogram of $\rho$ with a bin size of 0.01 when $\phi=0.4$. We can see from the figure that there is a small peak in the low-density range from $\rho=0.08$ to $0.09$. In the experiments, we set $\phi=0.4$ and removed all pixels $\bm{a}_{i}$ satisfying $\rho(i;\phi)\leq\omega$ for a truncation parameter $\omega$ specified in advance. We ran the same methods as in Section VI-B for Urban while varying $\omega$ from 00 to $0.15$.
@@
First, let us examine the experiments on Urban with $\phi=0.4$ and $\omega=0.1$. Under these settings, there were $2955$ pixels $\bm{a}_{i}$ ($3\%$ for total number of pixels) satisfying $\rho(i;\phi)\leq\omega$. We removed all of those pixels. For EEHT, we set the input parameters $\zeta$ and $\eta$ as $(\zeta,\eta)=(50,300)$. For MERIT, we used the same parameter values as in Section VI-B. We set $\lambda$ to $10^{-6},10^{-5},...,10^{3}$. For each $\lambda$ value, we ran MERIT and calculated the average MRSA values for six endmembers. The averages were at their minimum when $\lambda=10^{2}$. Therefore, we reported the results of MERIT with $\lambda=10^{2}$.
Table II summarizes the MRSA values for each endmember given in (10) and their averages given in (11). We can see from the table that the MRSA values of EEHT-C are smaller than those of the other methods, except in the case of Roof 2. Regarding computational time, EEHT-A, -B, and -C each took approximately 10 hours.
TABLE II: MRSA values (×$\times 10^{2}$) of methods for Urban with $\phi=0.4$ and $\omega=0.1$. The best scores are highlighted in bold.
@@
To assess the effectiveness of EEHT-C further, we computed abundance maps by using the estimated endmember signatures and compared them with the ground truth derived from the reference signatures. Figure 6 in Appendix F displays them. We can see that the abundance maps obtained by EEHT-C are similar to those obtained by the reference signatures except Roof 2.
@@
Next, let us examine the experiments on Urban with $\phi=0.4$ and varying $\omega$ from 00 to $0.15$ in $0.025$ increments. Figure 4 plots the average MRSA values. We can see from the figure that EEHT-C performs better than the other methods in average MRSA, except when $\omega=0.15$.
@@
The average MRSA values of EEHT-C reach their minimum at $\omega=0.1$. If $\omega$ is set to a large value, some of the pixels $\bm{a}_{i}$ satisfying $\rho(i;\phi)\leq\omega$ are no longer isolated and may be close to pure pixels. Removing such pixels can affect the number of pure pixels in an HSI. To explore this, we set $\phi=0.4$ and $\omega=0.125$ and removed all pixels $\bm{a}_{i}$ satisfying $\rho(i;\phi)\leq\omega$. At this setting, no pixels existed with an MRSA value of $0.05$ relative to the reference signature of Roof 1. Accordingly, when $\omega\geq 0.125$, it was impossible for EEHT-C to identify pixels containing the spectral signature of Roof 1 with high purity. For this reason, the average MRSA values of EEHT-C increased.
@@
As shown in Figure 4, the endmember extraction performance of the methods depends on the choice of $\omega$. Details of the observation for choosing the value of $\omega$ are provided in Appendix E.
@@
Figure 4: Average MRSA values (×$\times 10^{2}$) of methods for Urban with $\phi=0.4$ and varying $\omega$ from 00 to $0.15$ in $0.025$ increments.
VII Concluding Remarks

### v3

@@ VII-A Computational Efficiency
We chose ten equally spaced points between 0 and $1$ and set them as noise intensity levels $\nu$. Let $\nu_{1},\ldots,\nu_{10}$ denote these levels in increasing order, so that $0=\nu_{1}\leq\cdots\leq\nu_{10}=1$. Each dataset contained ten HSI matrices $A=W[I,\bar{H}]+V$ of size $50\times n$ with $r=10$ satisfying $\nu_{i}=\|V\|_{1}$ for each $i=1,\ldots,10$. We set $n$ from $500$ to $2500$ in $500$ increments and, thus, constructed five datasets in total.
@@
Figure 1: Average elapsed time (in seconds) of four methods for five datasets.
We compared RCE-SR and -DIR with CPLEX-SR and -DIR, since the Hottopixx methods in [17, 20, 31] use CPLEX and solve the Hottopixx models directly. Figure 1 summarizes the average elapsed time for each dataset. We can see from the figure that RCE is faster than CPLEX for $n\geq 1000$. In particular, RCE-SR is significantly faster than the other methods and its average elapsed time increases more slowly. RCE-DIR is faster than CPLEX-DIR, but its average elapsed time increases more rapidly than that of RCE-SR. The results suggest that RCE-SR should be used for implementing Hottopixx when dealing with large matrices.
To better understand why RCE-SR was faster than the other methods, we provide Tables I and II, which summarize the results for ten HSI matrices with varied noise intensity levels $\nu_{1},\ldots,\nu_{10}$ in the dataset with $n=2500$. For RCE-SR, Table II shows that $\ell$ did not increase significantly during the iterations, which in turn led to a substantial reduction in computation time compared with CPLEX-SR. For RCE-DIR, Table I indicates that the maximum values of $\ell$ exceeded $n/2$ for $\nu_{5},\ldots,\nu_{10}$ and in particular, that it solved $\mathsf{P}(\mathcal{L},\mathcal{L})$ and $\mathsf{D}(\mathcal{L},\mathcal{L})$ four times instead of three for $\nu_{5}$ and $\nu_{9}$. Consequently, the elapsed times of RCE-DIR were nearly the same as those of CPLEX-DIR for $\nu_{5}$ and $\nu_{9}$.
We also examined the effect of the parameters $\zeta$ and $\eta$ on the computational time of RCE-SR. In the experiments, $\zeta$ was varied from $5$ to $20$ in increments of $5$, and $\eta$ was varied from $50$ to $200$ in increments of $50$, resulting in $16$ parameter combinations. For each combination of $(\zeta,\eta)$, RCE-SR was applied to six datasets, comprising the five datasets used in the previous experiments and an additional dataset with $n=5000$. The results indicated that the elapsed time of RCE-SR did not change substantially when $(\zeta,\eta)$ were varied across the $16$ combinations. Detailed experimental results are provided in Appendix E.
TABLE I: Comparison of RCE-DIR and CPLEX-DIR on ten HSI matrices with noise intensity levels $\nu_{1},\ldots,\nu_{10}$ in the dataset with $n=2500$. The table reports the elapsed time (top block), the maximum value of $\ell$ observed during the iterations (middle block), and the number of times $\mathsf{P}(\mathcal{L},\mathcal{L})$ and $\mathsf{D}(\mathcal{L},\mathcal{L})$ were solved (bottom block). Entries correspond to $\nu_{1},\ldots,\nu_{10}$, ordered from left to right.
Elapsed time in seconds ($\times 10^{-2}$): RCE-DIR (top) and CPLEX-DIR (bottom)
1.2 6.5 8.9 50.1 110.9 173.0 97.4 85.4 99.9 38.6
69.0 444.8 174.4 119.5 121.5 214.2 169.7 158.7 92.1 64.3
Maximum value of $\ell$ during the iterations of RCE-DIR
200 424 709 995 1256 1340 1525 1614 1630 1487
Number of times RCE-DIR solved $\mathsf{P}(\mathcal{L},\mathcal{L})$ and $\mathsf{D}(\mathcal{L},\mathcal{L})$
1 4 3 3 4 3 3 3 4 3
TABLE II: Comparison of RCE-SR and CPLEX-SR on ten HSI matrices with noise intensity levels $\nu_{1},\ldots,\nu_{10}$ in the dataset with $n=2500$. The table reports the elapsed time (top block), the maximum value of $\ell$ observed during the iterations (middle block), and the number of times $\mathsf{P}(\mathcal{L},\mathcal{L})$ and $\mathsf{D}(\mathcal{L},\mathcal{L})$ were solved (bottom block). Entries correspond to $\nu_{1},\ldots,\nu_{10}$, ordered from left to right.
Elapsed time in seconds ($\times 10^{-2}$): RCE-SR (top) and CPLEX-SR (bottom)
1.1 3.4 3.4 3.4 5.8 4.2 3.0 3.6 5.1 3.2
31.9 109.9 175.8 52.5 50.2 48.2 38.0 39.4 36.5 44.0
Maximum value of $\ell$ during the iterations of RCE-SR
200 239 276 308 516 517 615 735 673 663
Number of times RCE-SR solved $\mathsf{P}(\mathcal{L},\mathcal{L})$ and $\mathsf{D}(\mathcal{L},\mathcal{L})$
1 3 3 3 4 3 2 2 3 2
VII-B Endmember Extraction Performance
We examined the endmember extraction performance of EEHT using synthetic HSI datasets. Experiments were first conducted on LMMs and subsequently on bilinear mixing models (BMMs) to assess the capability of EEHT in the presence of multiple photon interactions. BMMs have attracted attention as models for the nonlinear unmixing of hyperspectral images; see, for example, [37, 36, 24]. The datasets were derived from the Jasper Ridge and Samson HSI datasets, and subimages were extracted following the procedure described in Section V of [40]:
@@ VII-B Endmember Extraction Performance
Figure 2 displays RGB images of these datasets. The endmember signatures in these regions, previously identified in [40], were used as reference signatures for generating the synthetic datasets.
Figure 2: RGB images of Jasper Ridge (left) and Samson (right).
First, we evaluated the performance of EEHT on synthetic datasets generated from LMMs. Let $A^{\mathrm{real}}$ be the HSI matrix for a dataset, and $\bm{w}^{\mathrm{ref}}_{1},\ldots,\bm{w}^{\mathrm{ref}}_{r}$ be the reference signature vectors. The following procedure was applied:
@@
    Compute the optimal solution $X$ to problem (4) by letting $(A,\mathcal{I})=(A^{\mathrm{real}},\mathcal{J})$ in the problem. Set $X(:,\mathcal{J})=I$ and then $H=X$.
@@
We constructed two datasets using the $W,H$ and $V$ obtained above: dataset 1 from Jasper Ridge and dataset 2 from Samson. Each dataset contained $20$ HSI matrices that were generated as follows: choose $20$ equally spaced points between 0 and $1$ as noise intensity levels $\nu$; and construct an HSI matrix $A=WH+(\nu/\|V\|_{1})\cdot V$ for each $\nu$. The HSI matrices $A$ were of size $198\times 10000$ with $4$ endmembers for dataset 1 and of size $156\times 9025$ with $3$ endmembers for dataset 2. Note that $A=A^{\mathrm{real}}$ holds if $\nu=\|V\|_{1}$, where $\|V\|_{1}\approx 0.61$ for dataset 1 and $\|V\|_{1}\approx 0.15$ for dataset 2.
@@
Figure 3: MRSA scores ($\times 10^{2}$) of EEHT and existing methods for dataset 1 (top) and dataset 2 (bottom).
@@
For a fair comparison with EEHT, we included the SVD-based dimensionality reduction, described in Section VI, in MERIT: i.e., first compute the top-$r$ truncated SVD $A_{r}=U_{r}\Sigma_{r}V_{r}^{\top}$ of the input matrix $A$; then run MERIT on $A^{\prime}=\Sigma_{r}V_{r}^{\top}$.
We ran EEHT and the six existing methods on datasets 1 and 2, and evaluated their performance using MRSA scores. Regarding MERIT, it is necessary to fix the parameter $\lambda$ of problem (6) and the parameter $\mu$ of the function $\phi_{\mu}$ shown in (7). In particular, the choice of $\lambda$ affects its endmember extraction performance. We set $\mu$ to the default value of $10^{-5}$, as suggested by the authors of [33]. We then set $\lambda$ to $10^{-6},10^{-5},...,10^{3}$ and ran MERIT for each of these 10 values. For each $\lambda$ value, we calculated the mean MRSA score across all intensity levels $\nu$. For dataset 1 (resp., dataset 2), the mean score was minimized when $\lambda=10^{2}$ (resp., $10^{-1}$). Therefore, we report the results of MERIT with $\lambda=10^{2}$ for dataset 1 and $\lambda=10^{-1}$ for dataset 2 in Figure 3.
@@
Figure 3 plots the MRSA scores of the methods for datasets 1 and 2. We can see from the figure that, in most cases, EEHT-C achieved better MRSA scores than EEHT-A, -B, and the five greedy methods. This result suggests that using EEHT-C is preferable to EEHT-A and -B. EEHT-C achieved worse MRSA scores than MERIT for dataset 1, while it performed better than MERIT for dataset 2. For the $20$ HSI matrices included in each dataset, the computation times of EEHT-A, -B and -C were approximately 30 minutes on average.
@@
Next, we examined the performance of the methods tested in the LMM experiments, when the effect of multiple photon interactions was taken into account in datasets 1 and 2. This phenomenon is often observed when monitoring forested areas [10, 38]. In such areas, the interfaces between the canopy and the ground form a layered structure, and light can be reflected by multiple materials before reaching a hyperspectral camera.
We assumed that the HSI data follow the generalized bilinear model (GBM) described in [25]. In the GBM, second-order photon scattering is accounted for by incorporating an additional term into the LMM. That is, the observed spectral signature $\bm{a}_{j}$ at the $j$th pixel is expressed as
$\bm{a}_{j}=\sum_{i=1}^{r}h_{ij}\bm{w}_{i}+\sum_{p=1}^{r-1}\sum_{q=p+1}^{r}\xi_{pqj}h_{pj}h_{qj}\bm{w}_{p}\odot\bm{w}_{q}+\bm{v}_{j}$
where $0\leq\xi_{pqj}\leq 1$ and $\odot$ denotes the Hadamard product. For the sake of simplicity, let $\bm{v}_{j}^{\prime}$ denote the term $\sum_{p=1}^{r-1}\sum_{q=p+1}^{r}\xi_{pqj}h_{pj}h_{qj}\bm{w}_{p}\odot\bm{w}_{q}$. We can then write $A=[\bm{a}_{1},\ldots,\bm{a}_{n}]$ as
$A=WH+V^{\prime}+V$
where $V^{\prime}=[\bm{v}^{\prime}_{1},\ldots,\bm{v}^{\prime}_{n}]$. Note that $V^{\prime}$ is determined by $W,H$, and $\xi_{pqj}$.
We synthetically generated two datasets based on GBMs: dataset 3 from Jasper Ridge and dataset 4 from Samson. Each dataset contained HSI matrices defined by $A=WH+(\nu^{\prime}/\|V^{\prime}\|_{1})\cdot V^{\prime}+(\nu/\|V\|_{1})\cdot V$ where $\nu$ is the noise intensity and $\nu^{\prime}$ is the interaction intensity. Dataset 3 used $W$, $H$, and $V$ obtained in the LMM experiments from Jasper Ridge, while dataset 4 used those from Samson. To construct $V^{\prime}$, we used $W$ and $H$ and drew $\xi_{pqj}$ from a uniform distribution on the interval $[0,1]$. For each dataset, we set $\nu^{\prime}=0.2$ and varied $\nu$ from $0.2$ to $0.8$ in increments of $0.2$. For each $\nu$, we generated $20$ HSI matrices by changing $V^{\prime}$, resulting in $80$ HSI matrices in total per dataset.
For datasets 3 and 4, we ran the methods with the same parameter settings as in the LMM experiments. To avoid ambiguity, we specify that the MERIT parameter $\lambda$ was set to $10^{2}$ for dataset 3 and $10^{-1}$ for dataset 4. Table III summarizes the mean MRSA scores of EEHT and the six existing methods for these datasets. The results show that EEHT-C achieved better MRSA scores than the other methods for both datasets when $\nu\geq 0.4$.
TABLE III: Mean MRSA scores ($\times 10^{2}$) for each noise intensity level $\nu$, averaged over 20 randomly generated HSI matrices at each level. Results for dataset 3 are presented in the first block, and those for dataset 4 in the second. The best scores are highlighted in bold.
Dataset 3
$\nu=0.2$ $\nu=0.4$ $\nu=0.6$ $\nu=0.8$
EEHT-A 16.4 19.5 20.3 20.0
EEHT-B 14.1 16.7 18.2 20.1
EEHT-C 14.1 16.1 14.5 18.7
MERIT 18.1 17.1 19.1 20.1
SPA 16.7 17.9 18.9 19.7
PSAP 23.8 17.1 18.4 19.7
ER 15.2 17.1 18.6 19.7
SNPA 9.6 17.9 19.0 19.9
VCA 15.5 17.6 19.4 20.6
Dataset 4
$\nu=0.2$ $\nu=0.4$ $\nu=0.6$ $\nu=0.8$
EEHT-A 27.3 20.9 13.8 25.5
EEHT-B 3.5 19.2 13.8 16.2
EEHT-C 5.3 7.1 12.8 14.1
MERIT 8.5 13.4 27.9 17.2
SPA 7.0 10.7 14.2 17.0
PSPA 4.4 10.2 14.7 17.6
ER 3.5 10.7 14.7 17.5
SNPA 6.1 10.7 14.2 17.0
VCA 22.8 26.1 13.2 17.1
VII-C Hyperspectral Unmixing of the Urban HSI dataset
@@
Finally, we conducted an experimental study on hyperspectral unmixing of the Urban HSI dataset. The image was taken over Copperas Cove, Texas, USA by the HYDICE sensor. It consists of $307\times 307$ pixels with $210$ bands from 400 nm to 2500 nm. Following the procedure in [40, 39], we removed dirty bands from the image. The resulting image had 162 clean bands. The experimental studies in [40, 39] showed that there are mainly 44-66 endmembers in the image. Following the settings used in [21], our experiments set the number of endmembers to $6$: Asphalt, Grass, Tree, Roof 1, Roof 2 and Soil. These six endmember signatures were identified in [39]. We used them as reference signatures. Figure 4 displays the RGB image of the Urban dataset.
Figure 4: RGB image of Urban.
@@
We did not achieve satisfactory results for Urban by applying the methods that were tested on the synthetic HSI datasets described in Section VII-B. The estimated endmember signatures were not sufficiently close to the reference ones. Consequently, we conducted data-specific preprocessing aimed at enhancing the endmember extraction performance of these methods. The preprocessing was based on a reasonable assumption that there are a number of pixels close to each pure pixel in an HSI. Indeed, this assumption holds in the case of the Urban dataset: see Table IV that summarizes the number of pixels within an MRSA value of $0.05$ relative to the reference signatures of the endmembers. Here, we say that pixels are isolated if there are almost no pixels close to them. Under this assumption, even if we remove isolated pixels, there would be little effect on the number of pure pixels. It is possible that some of the isolated pixels contain large amounts of noise, which would worsen the endmember extraction performance of the methods.
@@
TABLE IV: Number of pixels within an MRSA value of $0.05$ relative to the reference signatures of the endmembers for Urban.
@@
To find isolated pixels, we introduced the neighborhood density of each pixel, defined as
@@
$\mathrm{NH}(i;\phi)=\left\{j\in\mathcal{N}\mid\ \mathrm{MRSA}(\bm{a}_{i},\bm{a}_{j})\leq\phi\right\},$
and the total number $n$ of pixels in the HSI. For Urban, we examined the neighborhood density $\rho$ while varying $\phi$. Figure 5 displays a histogram of $\rho$ with a bin size of 0.01 when $\phi=0.4$. We can see from the figure that there is a small peak in the low-density range from $\rho=0.08$ to $0.09$. In the experiments, we set $\phi=0.4$ and removed all pixels $\bm{a}_{i}$ satisfying $\rho(i;\phi)\leq\omega$ for a truncation parameter $\omega$ specified in advance. We ran the same methods as in Section VII-B for Urban while varying $\omega$ from 0 to $0.15$.
@@
First, let us examine the experiments on Urban with $\phi=0.4$ and $\omega=0.1$. Under these settings, there were $2955$ pixels $\bm{a}_{i}$ ($3\%$ of the total number of pixels) satisfying $\rho(i;\phi)\leq\omega$. We removed all of those pixels. For EEHT, we set the parameters $\zeta$ and $\eta$ to $(\zeta,\eta)=(50,300)$. For MERIT, we set the parameter $\lambda$ to $10^{-6},10^{-5},...,10^{3}$, while fixing the parameter $\mu$ to its default value of $10^{-5}$. For each $\lambda$ value, we ran MERIT and calculated the MRSA score. The score was minimized when $\lambda=10^{2}$. Therefore, we report the results of MERIT with $\lambda=10^{2}$.
Table V summarizes the MRSA values for each endmember and their averages (i.e., the MRSA scores). We can see from the table that the MRSA values of EEHT-C are smaller than those of the other methods, except in the case of Roof 2. Regarding computational time, EEHT-A, -B, and -C each took approximately 10 hours.
TABLE V: MRSA values ($\times 10^{2}$) of the methods for Urban with $\phi=0.4$ and $\omega=0.1$. The best scores are highlighted in bold.
@@
To assess the effectiveness of EEHT-C further, we computed abundance maps by solving problem (4) using the estimated endmember signatures, and compared them with the ground truth derived from the reference signatures. Figure 6 displays the resulting maps. We can see that the abundance maps obtained by EEHT-C are similar to those obtained by the reference signatures, except Roof 2. In addition, we quantitatively evaluated the accuracy of the abundance maps. Let $\hat{\bm{h}}_{1},\ldots,\hat{\bm{h}}_{n}\in\mathbb{R}^{r}$ be the abundance vectors estimated by the methods, and let $\bm{h}_{1},\ldots,\bm{h}_{n}\in\mathbb{R}^{r}$ be the corresponding ground truth vectors derived from the reference signatures. The root-mean-square error (RMSE) of the abundance fractions is defined as:
$\sqrt{\frac{1}{rn}\sum_{j=1}^{n}\|\hat{\bm{h}}_{j}-\bm{h}_{j}\|_{2}^{2}}.$
We also evaluated the reconstruction error (RE) of the unmixing process. To compute RE, we first construct the estimated endmember signature matrix $\hat{W}=[\bm{a}_{i_{1}},\ldots,\bm{a}_{i_{r}}]$ using the columns $\bm{a}_{i_{1}},\ldots,\bm{a}_{i_{r}}$ of the HSI matrix $A\in\mathbb{R}^{d\times n}$ returned by the methods. Each reconstructed column $\hat{\bm{a}}_{j}$ is then defined as $\hat{\bm{a}}_{j}=\hat{W}\hat{\bm{h}}_{j}$. Given the original columns $\bm{a}_{1},\ldots,\bm{a}_{n}$ of $A$ and their reconstructions $\hat{\bm{a}}_{1},\ldots,\hat{\bm{a}}_{n}$, RE is defined as:
$\sqrt{\frac{1}{dn}\sum_{j=1}^{n}\|\hat{\bm{a}}_{j}-\bm{a}_{j}\|_{2}^{2}}.$
Table VI summarizes the RMSE and RE values for each method, indicating that EEHT-C achieved the best RMSE and shared the best RE performance with VCA.
TABLE VI: RMSE ($\times 10^{2}$) and RE ($\times 10^{2}$) values of the abundance maps estimated by each method for Urban with $\phi=0.4$ and $\omega=0.1$. The best scores are highlighted in bold.
EEHT-C MERIT SPA PSPA ER VCA SNPA
RMSE 11.1 31.3 26.5 16.2 19.2 28.3 19.3
RE 14.6 25.3 32.7 28.5 33.9 14.6 32.8
@@
Next, let us examine the experiments on Urban with $\phi=0.4$ and varying $\omega$ from 0 to $0.15$ in increments of $0.025$. Figure 7 shows the MRSA scores. We can see from the figure that EEHT-C achieved better MRSA scores than the other methods, except when $\omega=0.15$.
@@
The MRSA scores of EEHT-C reached their minimum at $\omega=0.1$. When $\omega$ is set to a large value, some of the pixels $\bm{a}_{i}$ satisfying $\rho(i;\phi)\leq\omega$ are no longer isolated and may be close to pure pixels. Removing such pixels can affect the number of pure pixels in the HSI. To explore this, we set $\phi=0.4$ and $\omega=0.125$ and removed all pixels $\bm{a}_{i}$ satisfying $\rho(i;\phi)\leq\omega$. At this setting, no pixels existed with an MRSA value of $0.05$ relative to the reference signature of Roof 1. Accordingly, when $\omega\geq 0.125$, it was impossible for EEHT-C to identify pixels containing the spectral signature of Roof 1 with high purity. For this reason, the MRSA scores of EEHT-C increased.
@@
As shown in Figure 7, the endmember extraction performance of the methods depends on the choice of $\omega$. Details of the observation for choosing the value of $\omega$ are provided in Appendix F.
@@
Figure 7: MRSA scores ($\times 10^{2}$) of methods for Urban with $\phi=0.4$ and varying $\omega$ from 0 to $0.15$ in increments of $0.025$.
VIII Concluding Remarks